A convex surface has a uniform radius of curvature ‘5R’. A wheel of radius ‘R’ is rolling without slipping on it with its centre of mass moving with constant speed ‘v’. The acceleration of the point ‘P’ in contact with convex surface is

A)

 

b)

c)
 
d)
 
 
 
 

Grade:11

1 Answers

Khimraj
5 years ago
centrifugal force F = mv2/r
acceleration = F/m = v2/r = v2/5R
Hope it clears.......................................................................................